OpenStudy (anonymous):

The center of the circle with equation (x - 3)2 + (y + 2)2 = 4 has coordinates ___. Answer (3, -2) (3, 2) (-3, -2) (-3, 2)

OpenStudy (anonymous):

(3, -2)

OpenStudy (anonymous):

the..general eqn with center at (h,k) is (x-h)^2 +(y-k)^2=r^2...so r=radius!